OpenStudy (anonymous):

if the current is 6 amps and the voltage is 15 volts, which is the resistance ?

OpenStudy (anonymous):

Use ohms law V = IR V/I = R

OpenStudy (anonymous):

not helpful

OpenStudy (anonymous):

Put V = 15 volts I = 6 amp V/I = 15/6 = 5/2 R = 2.5 ohms